# Integral canonical models of exceptional Shimura varieties

### v1

@@ 4 Criteria for local integral canonical models
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{S}/\mathcal{O}$ be a smooth $\mathcal{O}$-scheme equipped with:
@@
    A strict $p$-torsion Fontaine-Laffaille module $\mathbf{M}\in\mathrm{MF}^{\nabla}_{[a,w+a],1}(\mathcal{S})$ whose Higgs bundle $\operatorname{gr}_{F}\mathbf{M}$ has quasifinite Kodaira–Spencer map.
Then if $p>2w$, $\mathcal{S}$ is an integral canonical model.
See Definition 2.2 and Definition 3.12 for the definitions of natural boundary and quasifinite Kodaira–Spencer map. In the normal case, we will additionally need a form of stratified resolution.
@@
    has natural boundary if each $\bar{\mathcal{S}}^{j}$ has natural boundary.
@@
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme equipped with:
@@
    A Higgs bundle $(V,\theta)$ on $\mathcal{Y}_{0}$ with quasifinite Kodaira–Spencer map.
@@
    For each $j$ a strict $p$-torsion Fontaine-Laffaille module $\mathbf{M}^{j}\in\mathrm{MF}^{\nabla}_{[a^{j},w^{j}+a^{j}],1}(\mathcal{S}^{j})$ and a nonzero torsion-free coherent sheaf $G^{j}$ on $\mathcal{S}^{j}$ such that
    $(G^{j},0)\otimes\operatorname{gr}_{F}\mathbf{M}^{j}$ $\cong(G^{j},0)\otimes\pi^{j*}(V,\theta)$
    $(G^{j},0)\otimes\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ $\cong(G^{j},0)\otimes\pi^{j*}(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$
    as Higgs sheaves.
Then if $p>2\max(w^{j})$, $\mathcal{Y}$ is an integral canonical model.
Note that Theorem 4.3 directly implies Theorem 4.1 by taking the trivial stratification, $(V,\theta)=\operatorname{gr}_{F}\mathbf{M}$, and $(G^{m},0)=(\mathcal{O}_{\mathcal{S}_{0}},0)$.
@@ 4.1 Extension in codimension $\geq 2$
We now use the results of Section 3.3 to show that, in the presence of a Fontaine–Laffaille module with quasi-finite Kodaira–Spencer map, morphisms from smooth sources which are defined outside of a set of codimension $\geq 2$ extend.
@@
Assume the setup of Theorem 4.3. Let $\mathcal{X}/\mathcal{O}$ be a smooth $\mathcal{O}$-scheme and $f:\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$ a morphism for $\mathcal{Z}\subset\mathcal{X}$ a subset of codimension $\geq 2$ which is contained in the special fiber. Then there is an extension $f:\mathcal{X}\to\mathcal{Y}$.
@@
If $j$ is minimal such that $Y^{j}$ contains the image of the generic fiber $X$, then the image of $X$ meets $Y^{j}\setminus Y^{j-1}$, so there is a dense open set of $X$ which lifts to $S^{j}$. By purity of the branch locus, $\mathcal{X}\setminus\mathcal{Z}\to\mathcal{X}$ induces an isomorphism on fundamental groups, so the map $X\to Y$ is $\mathcal{O}$-admissible. It follows by Lemma 2.3 that there is a finite type flat $\mathcal{X}^{\prime}/\mathcal{O}$, a proper morphism $\mathcal{X}^{\prime}\to\mathcal{X}$ which is an isomorphism on a dense open subset, and a morphism $\mathcal{X}^{\prime}\to\mathcal{S}^{j}$ whose generic fiber lifts the morphism $X\to Y^{j}$.
@@
Assume first that $\dim\mathcal{X}=2$. By resolution of singularities for surfaces ([lipman], specifically [stacks-project, Tag 0AHI]), there is a composition of blow-ups at closed points $\mathcal{X}_{n}\to\cdots\to\mathcal{X}_{1}\to\mathcal{X}$ whose composition factors as $\mathcal{X}_{n}\to\mathcal{X}^{\prime}\to\mathcal{X}$. To show that $\mathcal{X}^{\prime}\to\mathcal{S}^{j}\to\mathcal{Y}$ factors through $\mathcal{X}^{\prime}\to\mathcal{X}$ (and hence the existence of the required extension), it suffices (for instance by Lemma 2.16) to show the following:
@@
Note that by the quasifiniteness assumption on the Kodaira–Spencer map of $(V,\theta)$, a component $B$ of the special fiber of $\mathcal{X}_{n}$ is contracted if and only if the Higgs field of $(V,\theta)$ vanishes on the image of $B$ in $\mathcal{Y}_{0}$. Moreover, for any such $B$, the Higgs field of $\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ also vanishes on $B$, by assumption (3). The exceptional curves form a tree and the components $B$ which are not contracted form a sub-tree whose extremal components have $d_{B,\operatorname{\mathsf{End}}(\mathbf{M}^{j})}\leq 1$, so by Proposition 3.9 it suffices to show that $g^{*}\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ contains a sub-Higgs sheaf of slope larger than $w^{j}$, where $g:B\to\mathcal{S}^{j}_{0}$ is the map on the special fiber restricted to $B$.
[…]
$(G,0)\otimes g^{*}\operatorname{\mathsf{End}}(\mathbf{M}^{j})\cong(G,0)\otimes h^{*}(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)\cong(G,0)\otimes(\alpha\circ\gamma)^{*}\operatorname{\mathsf{End}}(\mathbf{M}^{i})$
where $G=g^{*}G^{j}\otimes(\gamma\circ\alpha)^{*}G^{i}$. Thus, $(G,0)\otimes\alpha^{*}(F^{*}T_{B^{\prime}},0)$ is a Higgs subsheaf of $(G,0)\otimes g^{*}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$. Since $B$ is rational, $G$ splits [^3] Possibly after a field extension. as a sum of factors $\mathcal{O}_{B}(d)$, and taking the largest degree factor we find that $((\alpha^{*}F^{*}T_{B^{\prime}})(e),0)$ is a Higgs subsheaf of $g^{*}\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M}^{j})$ for some $e\geq 0$. Since $\deg(\alpha^{*}F^{*}T_{B^{\prime}})(e)\geq 2p$, the conclusion follows. ∎
@@
[…]
@@ 4.2 Local models of images of period maps
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme equipped with:
@@
    A smooth $\mathcal{P}/\mathcal{O}$ with a proper dominant $\phi:\mathcal{P}\to\mathcal{Y}$ and a Fontaine-Laffaille module $\mathbf{M}\in\mathrm{MF}^{\nabla}_{[a,w+a]}(\mathcal{P})$ with mod $p$ reduction $\mathbf{M}_{0}$ such that:
@@
      The filtered flat vector bundle $(M_{0},\nabla,F^{\bullet}M_{0})$ underlying $\mathbf{M}_{0}$ is pulled back from $(V,\nabla,F^{\bullet}V)$ on $\mathcal{Y}_{0}$.
@@
      The Higgs bundle $(\operatorname{gr}_{F}V,\operatorname{gr}_{F}\nabla)$ has quasifinite Kodaira–Spencer map.
@@
      The local system ${}_{et}\mathbb{V}_{p}$ on $P$ corresponding to $\mathbf{M}$ under the Fontaine–Laffaille correspondence descends to $Y$.
@@
    ${\mathcal{T}^{j}}$${\mathcal{P}}$${\mathcal{S}^{j}}$${\mathcal{Y}^{\prime j}}$${\mathcal{Y}}$$\scriptstyle{q^{j}}$$\scriptstyle{t^{j}}$$\scriptstyle{h}$$\scriptstyle{\pi^{j}}$
@@
Then if $p>2w$, $\mathcal{Y}$ is an integral canonical model.
@@
Let $F$ be a mixed characteristic local field with ring of integers $\mathcal{O}$ and residue field $k$ of characteristic $p$. Let $\mathcal{S},\mathcal{T}/\mathcal{O}$ be smooth.
@@
    If $f:\mathcal{T}\to\mathcal{S}$ is finite étale of prime to $p$ degree and ${}_{et}\mathbb{V}_{p}$ is a $\mathbb{Z}_{p}$-local system on $S$ such that $f^{*}_{et}\mathbb{V}_{p}$ is crystalline, then ${}_{et}\mathbb{V}_{p}$ is crystalline, and the inclusion ${}_{et}\mathbb{V}_{p}\to f_{*}f^{*}(_{et}\mathbb{V}_{p})$ respects the lift of Frobenius.
@@
      $f:\mathcal{T}\to\mathcal{S}$ a proper dominant morphism, with base-change $f_{0}:\mathcal{T}_{0}\to\mathcal{S}_{0}$ to the special fiber;
@@
      $\mathcal{U}_{0}\subset\mathcal{S}_{0}$ a dense open set with preimage $\mathcal{V}_{0}=f_{0}^{-1}(\mathcal{U}_{0})$;
@@
      strict $p$-torsion Fontaine–Laffaille modules $\mathbf{M}=(M,\nabla,F^{\bullet}M,\Phi_{\mathbf{M}})$ on $\mathcal{T}$ and $\mathbf{N}_{\mathcal{U}}=(N_{\mathcal{U}},\nabla,F^{\bullet}N_{\mathcal{U}},\Phi_{\mathbf{N}_{\mathcal{U}}})$ on $\mathcal{U}$ as well as an extension $(N,\nabla,F^{\bullet}N)$ of the underlying filtered flat vector bundle of $\mathbf{N}_{\mathcal{U}}$ to $\mathcal{S}_{0}$;
@@
      compatible isomorphisms $f^{*}\mathbf{N}_{\mathcal{U}}\xrightarrow{\cong}\mathbf{M}|_{\mathcal{V}}$ and $f_{0}^{*}(N,\nabla,F^{\bullet}N)\xrightarrow{\cong}(M,\nabla,F^{\bullet}M)$.
    Then there is a $p$-torsion Fontaine–Laffaille module $\mathbf{N}$ extending $\mathbf{N}_{\mathcal{U}}$ and an isomorphism $f_{0}^{*}\mathbf{N}\to\mathbf{M}$ compatible with the above isomorphisms.
@@
The category of Fontaine–Laffaille modules is a full subcategory of the category of local systems, so the descent datum of $f^{*}_{et}\mathbb{V}_{p}$ also descends the corresponding Fontaine–Laffaille modules [faltings]. Since ${}_{et}\mathbb{V}_{p}\to f_{*}f^{*}(_{et}\mathbb{V}_{p})$ is a direct factor of local systems, it follows it comes from a direct factor of Fontaine–Laffaille modules as well.
We now prove (2). Observe that the $p$-curvature of $(N,\nabla)$ is nilpotent of the same index as $\mathbf{N}$ (and $\mathbf{M}$). Unwrapping the definitions, we have isomorphisms
$\Phi_{\mathbf{N}_{\mathcal{U}}}:C_{\mathcal{U}}^{-1}\operatorname{gr}_{F}\mathbf{N}|_{\mathcal{U}_{0}}\xrightarrow{\cong}(N,\nabla)|_{\mathcal{U}_{0}}$
$\Phi_{\mathbf{M}}:C_{\mathcal{T}}^{-1}\operatorname{gr}_{F}\mathbf{M}\xrightarrow{\cong}(M,\nabla)$
such that $f_{0}^{*}\Phi_{\mathbf{N}_{\mathcal{U}}}=\Phi_{\mathbf{M}}$ over $\mathcal{V}_{0}$, using the natural identification $f^{*}_{0}C^{-1}_{\mathcal{S}}\operatorname{gr}_{F}\mathbf{N}\cong C^{-1}_{\mathcal{T}}f_{0}^{*}\operatorname{gr}_{F}\mathbf{N}\cong C^{-1}_{\mathcal{T}}\operatorname{gr}_{F}\mathbf{M}$. Thus, we have solid commutative diagrams
@@
Since $\mathcal{T},\mathcal{S}$ are smooth and $f$ is proper dominant (and therefore the same for $f_{0}$), $\mathcal{O}_{\mathcal{S}_{0}}\to f_{0*}\mathcal{O}_{\mathcal{T}_{0}}$ has torsion-free cokernel. As $N$ and $C^{-1}_{\mathcal{S}}\operatorname{gr}_{F}\mathbf{N}$ are locally free, the vertical arrows in the left square of (4.2.1) have torsion-free cokernel as well, and it follows that $f_{0*}\Phi_{\mathbf{M}}$ restricts to an extension $\Phi_{\mathbf{N}}:C^{-1}_{\mathcal{S}}\operatorname{gr}_{F}\mathbf{N}\xrightarrow{\cong}(N,\nabla)$ of $\Phi_{\mathbf{N}_{\mathcal{U}}}$. ∎
@@
By assumption, for each $j$, we have a Fontaine–Laffaille module $t^{j*}\mathbf{M}$ on $\mathcal{T}^{j}$ whose corresponding local system is pulled back from $\mathcal{Y}$. Moreover, the underlying filtered flat vector bundle of $t^{j*}\mathbf{M}_{0}$ is pulled back from $(V,\nabla,F^{\bullet}V)$ on $\mathcal{Y}_{0}$. Applying part 1 of Lemma 4.8 to the open set $\mathcal{U}^{j}$ of $\mathcal{S}^{j}$ over which $q^{j}$ is finite étale, we conclude that there is a Fontaine–Laffaille module $\mathbf{N}_{\mathcal{U}^{j}}$ which pulls back to $(t^{j*}\mathbf{M})|_{(q^{j})^{-1}(\mathcal{U}^{j})}$. Applying part 2, we find that $(\mathbf{N}_{\mathcal{U}^{j}})_{0}$ extends to a Fontaine–Laffaille module $\mathbf{N}^{j}_{0}$ which pulls back to $t^{j*}\mathbf{M}_{0}$. In particular, we have $q^{j*}\operatorname{gr}_{F}\mathbf{N}^{j}_{0}\cong q^{j*}\pi^{j*}(\operatorname{gr}_{F}V,\operatorname{gr}_{F}\nabla)$, so
$(q^{j}_{*}\mathcal{O}_{\mathcal{T}^{j}_{0}},0)\otimes\operatorname{gr}_{F}\mathbf{N}^{j}_{0}\cong(q^{j}_{*}\mathcal{O}_{\mathcal{T}_{0}^{j}},0)\otimes\pi^{j*}(\operatorname{gr}_{F}V,\operatorname{gr}_{F}\nabla)$
as Higgs sheaves. By the same argument, we have
$(q^{j}_{*}\mathcal{O}_{\mathcal{T}^{j}_{0}},0)\otimes\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{N}^{j})\cong(q^{j}_{*}\mathcal{O}_{\mathcal{T}_{0}^{j}},0)\otimes\pi^{j*}(\operatorname{gr}_{F}\operatorname{\mathsf{End}}(V),\operatorname{gr}_{F}\operatorname{ad}\nabla)$
as Higgs sheaves. Thus, all the conditions of Theorem 4.3 are satisfied. ∎

### v2

@@ 4 Criteria for local integral canonical models
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension with ring of integers $\mathcal{O}$. Let $\mathcal{S}/\mathcal{O}$ be a smooth $\mathcal{O}$-scheme equipped with:
@@
    A Fontaine-Laffaille module $\mathbf{M}\in\mathrm{MF}_{[0,w]}(\hat{\mathcal{S}})$ with ample Griffiths bundle.
Then if $p>w+1$, $\mathcal{S}$ is an integral canonical model.
See Definition 2.2 and Definition 2.27 for the definitions of ($\pi$-)natural boundary and the Griffiths bundle. In the normal case, we will additionally need a form of stratified resolution.
@@
    has natural boundary if each $\bar{\mathcal{S}}^{j}$ has $\pi^{j}$-natural boundary.
@@
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension with ring of integers $\mathcal{O}$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme equipped with:
@@
    A filtered bundle $(V,F^{\bullet}V)$ on $\hat{\mathcal{Y}}$ with ample Griffiths bundle.
@@
    For each $j$ a Fontaine-Laffaille module $\mathbf{M}^{j}\in\mathrm{MF}_{[0,w^{j}]}(\hat{\mathcal{S}}^{j})$ and a proper dominant generically finite morphism $q^{j}:\mathcal{T}^{j}\to\mathcal{S}^{j}$ from a smooth $\mathcal{T}^{j}/\mathcal{O}$ such that
    $\hat{q}^{j*}(M^{j},F^{\bullet}M^{j})$ $\cong\hat{q}^{j*}\hat{\pi}^{j*}(V,F^{\bullet}V)$
    as filtered vector bundles.
Then if $p>\max(w^{j})+1$, $\mathcal{Y}$ is an integral canonical model.
Note that Theorem 4.3 directly implies Theorem 4.1 by taking the trivial stratification and $\mathcal{T}^{m}=\mathcal{S}^{m}=\mathcal{Y}=\mathcal{S}$.
@@ 4.1 Extension in codimension $\geq 2$
We now use the results of Section 3 to show that, in the presence of a Fontaine–Laffaille module with ample Griffiths bundle, morphisms from smooth sources which are defined outside of a set of codimension $\geq 2$ extend.
@@
Assume the setup of Theorem 4.3. Let $\mathcal{X}/\mathcal{O}$ be a smooth finite type $\mathcal{O}$-scheme and $f:\mathcal{X}\setminus\mathcal{Z}\to\mathcal{Y}$ a morphism for $\mathcal{Z}\subset\mathcal{X}$ a subset of codimension $\geq 2$ which is contained in the special fiber. Then there is an extension $f:\mathcal{X}\to\mathcal{Y}$.
@@
If $j$ is minimal such that $Y^{j}$ contains the image of the generic fiber $X$, then the image of $X$ meets $Y^{j}\setminus Y^{j-1}$, so there is a dense open set of $X$ which lifts to $S^{j}$. By purity of the branch locus, $\mathcal{X}\setminus\mathcal{Z}\to\mathcal{X}$ induces an isomorphism on fundamental groups, so the map $X\to Y$ is $\mathcal{O}$-admissible. It follows by Lemma 2.4 that there is a finite type flat $\mathcal{X}^{\prime}/\mathcal{O}$, a proper morphism $\pi:\mathcal{X}^{\prime}\to\mathcal{X}$ which is an isomorphism on a dense open subset, and a morphism $\mathcal{X}^{\prime}\to\mathcal{S}^{j}$ whose generic fiber lifts the morphism $X\to Y^{j}$.
@@
Assume first that $\dim\mathcal{X}=2$, so there is a finite set of points $\Sigma$ in the special fiber of $\mathcal{X}$ such that $\pi$ is an isomorphism on the complement $\mathcal{X}\setminus\Sigma$. By resolution of singularities for surfaces ([lipman], specifically [stacks-project, Tag 0AHI]), there is a composition of blow-ups at points lying above $\Sigma$, $\mathcal{X}_{n}\to\cdots\to\mathcal{X}_{1}\to\mathcal{X}$, whose composition factors as $\mathcal{X}_{n}\to\mathcal{X}^{\prime}\to\mathcal{X}$. Note that the map $\mathcal{X}\setminus\Sigma\to\mathcal{Y}$ lifts to $\mathcal{S}^{j}$. To show that $\mathcal{X}^{\prime}\to\mathcal{S}^{j}\to\mathcal{Y}$ factors through $\mathcal{X}^{\prime}\to\mathcal{X}$ (and hence the existence of the required extension), it suffices (for instance by Lemma 2.17) to show the following:
@@
Applying Theorem 3.1 to the diagram
${\mathcal{X}_{n}}$${\mathcal{S}^{j}}$${\mathcal{X}\setminus\Sigma}$${\mathcal{X}}$
we obtain that the Griffiths bundle of $\mathbf{M}^{j}$ is trivial on every exceptional divisor $E$. For each $E$ we may take a curve $C$ in $\mathcal{T}^{j}$ dominating $E$, and it follows that the Griffiths bundle of $V$ is trivial on $C$, hence it is degree 0 on $E$. Thus, every exceptional divisor is contracted in $\mathcal{Y}$. ∎
@@
The $\dim\mathcal{X}=2$ step in the proof of Proposition 4.4 in the setting of Theorem 4.1 follows immediately from the last claim of Theorem 3.1.
@@ 4.2 Local models of images of period maps
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension with ring of integers $\mathcal{O}$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme equipped with:
@@
    A smooth $\mathcal{P}/\mathcal{O}$ with a proper dominant $h:\mathcal{P}\to\mathcal{Y}$ with $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}h_{*}\mathcal{O}_{\mathcal{P}}$ and a Fontaine-Laffaille module $\mathbf{M}\in\mathrm{MF}_{[0,w]}(\hat{\mathcal{P}})$ such that:
@@
      The filtered flat vector bundle $(M,\nabla,F^{\bullet}M)$ underlying $\mathbf{M}$ is pulled back from $(V,\nabla,F^{\bullet}V)$ on $\hat{\mathcal{Y}}$.
@@
      The Griffiths bundle of $(V,\nabla,F^{\bullet}V)$ is ample.
@@
      The local system ${}_{et}\mathbb{V}_{p}$ on $\mathcal{P}^{rig}$ corresponding to $\mathbf{M}$ under the Fontaine–Laffaille correspondence descends to $\mathcal{Y}^{rig}$.
@@
    ${\mathcal{T}^{j}}$${\mathcal{P}}$${\mathcal{S}^{j}}$${\mathcal{Y}^{\prime j}}$${\mathcal{Y}}$$\scriptstyle{q^{j}}$$\scriptstyle{t^{j}}$$\scriptstyle{h}$$\scriptstyle{\pi^{\prime j}}$
@@
Then if $p>w+1$, $\mathcal{Y}$ is an integral canonical model.
@@
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension with ring of integers $\mathcal{O}$. Let $\mathcal{S},\mathcal{T}$ be smooth $\mathcal{O}$-schemes.
@@
    If $f:\mathcal{T}\to\mathcal{S}$ is finite étale of prime to $p$ degree and ${}_{et}\mathbb{V}_{p}$ is a $\mathbb{Z}_{p}$-local system on $\mathcal{S}^{rig}$ such that $(f^{rig})^{*}(_{et}\mathbb{V}_{p})$ is crystalline, then ${}_{et}\mathbb{V}_{p}$ is crystalline.
@@
      $f:\mathcal{T}\to\mathcal{S}$ a proper dominant morphism;
@@
      $\mathcal{U}\subset\mathcal{S}$ a dense open set with preimage $\mathcal{V}=f^{-1}(\mathcal{U})$ such that $f|_{\mathcal{V}}:\mathcal{V}\to\mathcal{U}$ is étale and $\mathcal{S}\setminus\mathcal{U}$ is flat over $\mathcal{O}$;
@@
      Fontaine–Laffaille modules $\mathbf{M}\in\operatorname{MF}_{[0,w]}(\hat{\mathcal{T}})$ and $\mathbf{N}_{\mathcal{U}}\in\operatorname{MF}_{[0,w]}(\hat{\mathcal{U}})$ as well as an extension $(N,\nabla,F^{\bullet}N)$ of the underlying filtered flat vector bundle of $\mathbf{N}_{\hat{\mathcal{U}}}$ to $\hat{\mathcal{S}}$;
@@
      an isomorphism of filtered flat bundles $\hat{f}^{*}(N,\nabla,F^{\bullet}N)\to(M,\nabla,F^{\bullet}M)$ whose restriction to $\mathcal{V}$ underlies an isomorphism $\hat{f}^{*}\mathbf{N}_{\hat{\mathcal{U}}}\xrightarrow{\cong}\mathbf{M}|_{\hat{\mathcal{V}}}$.
    Then there is a Fontaine–Laffaille module $\mathbf{N}$ on $\hat{\mathcal{S}}$ and an isomorphism $\hat{f}^{*}\mathbf{N}\to\mathbf{M}$ extending the above isomorphism.
@@
The category of Fontaine–Laffaille modules is a full subcategory of the category of local systems, so the descent datum of $f^{*}_{et}\mathbb{V}_{p}$ also descends the corresponding Fontaine–Laffaille modules [faltings].
We now prove (2). The Frobenius morphism on $\mathbf{M}$ gives a morphism $\phi_{M}:\hat{f}^{*}F^{*}(M,\nabla)\to\hat{f}^{*}(M,\nabla)$ which is compatible with the one on $\hat{f}^{*}(N,\nabla)$, and we have natural solid commutative diagrams
@@
where in the top row we push-forward just the underlying vector bundles. Since $\mathcal{T},\mathcal{S}$ are smooth and $f$ is proper dominant, $\mathcal{O}_{\mathcal{S}}\to f_{*}\mathcal{O}_{\mathcal{T}}$ has torsion-free cokernel. Indeed, since $\mathcal{S},\mathcal{T}$ are smooth, $\mathcal{O}_{\mathcal{S}}\to f_{*}\mathcal{O}_{\mathcal{T}}$ is a split summand outside codimension 2, and both sheaves are reflexive. As $N$ and $F^{*}N$ are locally free, the vertical arrows in the left square of (4.2.1) have torsion-free cokernel as well, and it follows that $\hat{f}_{*}\phi_{M}$ restricts to an extension $\phi:F^{*}(N,\nabla)\xrightarrow{\cong}(N,\nabla)$. The extension $\phi$ is flat since it is on $\hat{\mathcal{U}}$.
For the strong divisibility condition, note that by eq. 3.1.1 $F^{*}(M,\nabla)$ has a natural flat filtration which is locally given by $\tilde{F}^{i}_{M}=\sum_{0\leq j\leq i}p^{j}\phi^{*}F^{i-j}$, and $\phi_{M}$ is divisible by $p^{i}$ on $\tilde{F}^{i}$. The strong divisibility condition is then equivalent to $p^{-w}\phi_{M}:\tilde{F}^{w}_{M}\to M$ being an isomorphism. Using the analog of the diagram eq. 4.2.1 for this morphism, since $p^{-w}\hat{f}_{*}\phi_{M}:\tilde{F}^{w}_{M}\to f_{*}M$ is an isomorphism and restricts to $p^{-w}\phi_{N}$ on $\hat{\mathcal{U}}$, we conclude that $p^{-1}w\phi_{N}$ is an isomorphism. ∎
@@
By assumption, for each $j$, we have a Fontaine–Laffaille module $\hat{t}^{j*}\mathbf{M}$ on $\hat{\mathcal{T}}^{j}$ whose corresponding local system is pulled back from $\mathcal{Y}^{rig}$. Moreover, the underlying filtered flat vector bundle of $\hat{t}^{j*}\mathbf{M}$ is pulled back from $(V,\nabla,F^{\bullet}V)$ on $\hat{\mathcal{Y}}$. Applying part 1 of Lemma 4.8 to the open set $\mathcal{U}^{j}$ of $\mathcal{S}^{j}$ over which $q^{j}$ is finite étale, we conclude that there is a Fontaine–Laffaille module $\mathbf{N}_{\hat{\mathcal{U}}^{j}}$ which pulls back to $(\hat{t}^{j*}\mathbf{M})|_{(\hat{q}^{j})^{-1}(\hat{\mathcal{U}}^{j})}$. Applying part 2, we find that $\mathbf{N}_{\hat{\mathcal{U}}^{j}}$ extends to a Fontaine–Laffaille module $\mathbf{N}^{j}$ which pulls back to $\hat{t}^{j*}\mathbf{M}$. In particular, we have $\hat{q}^{j*}(N^{j},F^{\bullet}N^{j})\cong\hat{q}^{j*}\hat{\pi}^{j*}(V,F^{\bullet}V)$. Thus, all the conditions of Theorem 4.3 are satisfied. ∎